# MAIN: Mutual Alignment Is Necessary for instruction tuning

### v2

@@ Abstract
Instruction tuning has empowered large language models (LLMs) to achieve remarkable performance, yet its success heavily depends on the availability of large-scale, high-quality instruction-response pairs. To meet this demand, various methods have been developed to synthesize data at scale. However, current methods for scaling up data generation often overlook a crucial aspect: the alignment between instructions and responses. We hypothesize that the quality of instruction-response pairs is determined not by the individual quality of each component, but by the degree of mutual alignment. To address this, we propose a Mutual Alignment Framework (MAIN) which enforces coherence between instructions and responses through mutual constraints. We demonstrate that MAIN generalizes well across model architectures and sizes, achieving state-of-the-art performance on LLaMA, Mistral, and Qwen models across diverse benchmarks. This work underscores the critical role of instruction-response alignment in enabling generalizable and high-quality instruction tuning for LLMs.
@@
Fanyi Yang1thanks: Work done during internship at Microsoft., Jianfeng Liu2, Xin Zhang2, Haoyu Liu2, Xixin Cao1, Yuefeng Zhan2, Hao Sun2, Weiwei Deng2, Feng Sun2, Qi Zhang2 1Peking University  2Microsoft Corporation yangfanyi@stu.pku.edu.cn, cxx@ss.pku.edu.cn {jianfengliu, xinzhang3, yuefzh, hasun, dedeng, sunfeng, qizhang}@microsoft.com implhy@gmail.com
@@ 2.3 Mutual Alignment
0:  Seed data $\mathcal{D}_{\text{seed}}=\{(I,R)\}$, Unlabeled data $\mathcal{D}_{\text{unlabeled}}=\{R_{u}\}$, Base model $M_{\text{base}}$, Number of iterations $N$
@@
6:     Generate pseudo-responses $\hat{R}$ from $I$ in $\mathcal{D}_{\text{seed}}$ using $M_{f}^{k+1}$
@@ 3.3 Baselines
AlpacaEval IFEval ARC_C MMLU HellaSwag Winogrande TruthfulQA
P-S I-S P-L I-L
@@ 4.1 Quantitative Results
We conduct experiments across three benchmarks, each assessing a different aspect of the MAIN: AlpacaEval evaluates output preference, IFEval measures instruction-following ability, and OpenLLM tests reasoning capability, with additional experiments detailed in Appendix B
@@ 4.3 Data Alignment Analysis
We assessed the alignment quality of instructions generated by our MAIN method compared to baselines using blind pairwise evaluations conducted by GPT-4. Specifically, we randomly selected 1000 responses from the Falcon RefinedWeb dataset. For each response, GPT-4 evaluated two candidate instructions—one from MAIN and one from a baseline—in random order to avoid positional bias. We then calculated Win, Tie, and Loss rates based on GPT-4’s judgments: Win indicates GPT-4 preferred MAIN, Tie indicates no clear preference, and Loss indicates GPT-4 preferred the baseline. Using the Qwen2.5-14B as the base model, MAIN consistently outperformed all baselines, with win rates ranging from 61.7% to 81.6%, demonstrating superior instruction alignment capability (see Table 3). The evaluation prompt used by GPT-4 is detailed in Appendix F.
@@ Appendix C Ablations
Table 7: Ablation study on the effect of iteration count $N$. We analyze the influence of varying the number of training iterations ($N=1,2,3,4,5,10,15,20$) on Llama-2-7B fine-tuned on Falcon-RefinedWeb.
@@
Appendix D Detailed Benchmark Settings
@@
Appendix E Computational Cost Analysis
@@
Appendix F GPT-4 Evaluation Prompt

### v3

@@ Abstract
Instruction tuning has empowered large language models (LLMs) to achieve remarkable performance, yet its success heavily depends on the availability of large-scale, high-quality instruction-response pairs. To meet this demand, various methods have been developed to synthesize data at scale. However, current methods for scaling up data generation often overlook a crucial aspect: the alignment between instructions and responses. We hypothesize that the quality of instruction-response pairs is determined not by the individual quality of each component, but by the degree of mutual alignment. To address this, we propose a Mutual Alignment Framework (MAIN) which enforces coherence between instructions and responses through mutual constraints. We demonstrate that MAIN generalizes well across model architectures and sizes, achieving state-of-the-art performance on LLaMA, Mistral, and Qwen models across diverse benchmarks. This work underscores the critical role of instruction-response alignment in enabling generalizable and high-quality instruction tuning for LLMs. All code is available from our repository.
@@
Fanyi Yang1thanks: Work done during internship at Microsoft., Jianfeng Liu2thanks: Corresponding Author., Xin Zhang2, Haoyu Liu2, Xixin Cao1, Yuefeng Zhan2, Hao Sun2, Weiwei Deng2, Feng Sun2, Qi Zhang2 1Peking University  2Microsoft Corporation yangfanyi@stu.pku.edu.cn, cxx@ss.pku.edu.cn {jianfengliu, xinzhang3, yuefzh, hasun, dedeng, sunfeng, qizhang}@microsoft.com implhy@gmail.com
@@ 2.3 Mutual Alignment
0: Seed data $\mathcal{D}_{\text{seed}}=\{(I,R)\}$, Unlabeled data $\mathcal{D}_{\text{unlabeled}}=\{R_{u}\}$, Base model $M_{\text{base}}$, Number of iterations NN
@@
6:  Generate pseudo-responses $\hat{R}$ from II in $\mathcal{D}_{\text{seed}}$ using $M_{f}^{k+1}$
@@ 3.3 Baselines
AlpacaEval
IFEval
ARC_C
MMLU
HellaSwag
Winogrande
TruthfulQA
P-S
I-S
P-L
I-L
@@ 4.1 Quantitative Results
We conduct experiments across three benchmarks, each assessing a different aspect of the MAIN: AlpacaEval evaluates output preference, IFEval measures instruction-following ability, and OpenLLM tests reasoning capability, with additional experiments detailed in Appendix B. For completeness, we also provide additional results on multilingual generalization in Appendix D.
@@ 4.3 Data Alignment Analysis
We assessed the alignment quality of instructions generated by our MAIN method compared to baselines using blind pairwise evaluations conducted by GPT-4. Specifically, we randomly selected 1000 responses from the Falcon RefinedWeb dataset. For each response, GPT-4 evaluated two candidate instructions—one from MAIN and one from a baseline—in random order to avoid positional bias. We then calculated Win, Tie, and Loss rates based on GPT-4’s judgments: Win indicates GPT-4 preferred MAIN, Tie indicates no clear preference, and Loss indicates GPT-4 preferred the baseline. Using the Qwen2.5-14B as the base model, MAIN consistently outperformed all baselines, with win rates ranging from 61.7% to 81.6%, demonstrating superior instruction alignment capability (see Table 3). The evaluation prompt used by GPT-4 is detailed in Appendix H.
@@ Appendix C Ablations
Table 7: Ablation study on the effect of iteration count NN. We analyze the influence of varying the number of training iterations ($N=1,2,3,4,5,10,15,20$) on Llama-2-7B fine-tuned on Falcon-RefinedWeb.
@@
Appendix D Multilingual Generalization
To evaluate multilingual robustness of MAIN, we perform experiments on diverse multilingual datasets.
Setup.
The backbone model is Qwen-2.5-14B-Base. Seed data contains 3.2k instruction–response pairs from the multilingual portion of OpenAssistant, evenly sampled across French (fr), Japanese (ja), Spanish (es), Arabic (ar), and Chinese (zh). Additionally, 500k responses are drawn from the multilingual mC4 corpus, a naturally web collection. After applying mutual alignment and filtering, we obtain a balanced set of 20k pairs for supervised fine-tuning. Baselines use identical data and training settings to ensure comparability.
Benchmarks.
Models are evaluated on multilingual reasoning and instruction-following benchmarks:
  • JMMLU (ja), AMMLU (ar), CMMLU (zh): 5-shot reasoning dataset.
  • mIFEval (fr/ja/es): multilingual instruction-following dataset.
Results.
Table 8 summarizes multilingual evaluation on instruction-following and reasoning benchmarks . MAIN attains the best performance on these benchmarks, with the largest gains on mIFEval (fr/ja/es) and notable improvements on JMMLU and CMMLU; performance on AMMLU is on par with the strongest baseline. These results demonstrate that our framework exhibits strong generalization ability across diverse multilingual settings.
Method mIFEval Reasoning
fr ja es JMMLU AMMLU CMMLU
Humpback 63.3 46.0 42.0 60.5 57.3 81.9
Better Align 65.3 50.1 43.2 62.7 59.8 80.6
MAIN 69.7 51.8 47.1 64.9 59.6 83.5
Table 8: Evaluation of multilingual instruction-following and reasoning dataset.
Appendix E Detailed Benchmark Settings
@@
Appendix F Mutual Filtering on MMLU
We compare LLaMA-2-7B trained with mutual alignment only (No-filter) and with mutual alignment plus mutual filtering (Filtered); all other settings are identical. Table 10 reports accuracy (%) by MMLU super-category.
MMLU Super-Category No-filter Filtered $\Delta$
STEM 42.7 45.5 +2.8
Humanities 45.1 46.3 +1.2
Social Sciences 46.7 47.2 +0.5
Other Professions 41.9 42.6 +0.7
Overall 44.1 45.4 +1.3
Table 10: MMLU accuracy (%) by super-category for LLaMA-2-7B. Both settings include mutual alignment; Filtered additionally applies mutual filtering when selecting pseudo-labeled pairs.
The largest gains arise in STEM, where instructions and responses tend to be tightly coupled and logically recoverable; such pairs pass bidirectional consistency checks at higher rates, yielding larger improvements. In more open-ended areas (e.g., Humanities, Social Sciences), valid responses are more diverse, so exact reconstruction is harder and gains are accordingly smaller. Overall, these deltas indicate that mutual filtering is most beneficial for categories with higher instruction–response determinism, while still providing modest positive effects elsewhere.
Appendix G Computational Cost Analysis
@@
Appendix H GPT-4 Evaluation Prompt
@@
Response:
{response}
Instruction A:
{instruction_A}
Instruction B:
{instruction_B}